# Timely Communications for Remote Inference

### v1

@@ IV-A Reaction Prediction with Delay
To facilitate understanding of the counter-intuitive experimental results illustrated in Figs. 2-3, we present the following analytical example for reaction prediction.
@@
Consider a causal system represented by $Y_{t}=f(X_{t-d})$, where $X_{t}$ and $Y_{t}$ are the input and output of the system, respectively, $d\geq 0$ is the delay introduced by the system, and $f(\cdot)$ is a function.
@@
If $X_{t}$ is a Markov chain and $Y_{t}=f(X_{t-d})$, then $H_{L}(Y_{t}|X_{t-\delta})$ decreases with $\delta$ when $0\leq\delta\leq d$ and increases with $\delta$ when $\delta\geq d$. In addition, for any random variable $Z$,
$H_{L}(Y_{t}|X_{t-d})=H_{L}(Y_{t}|Y_{t})\leq H_{L}(Y_{t}|Z).$ (29)
@@
Lemma 4 implies that the feature $X_{t-d}$ achieves the minimum expected loss in predicting $Y_{t}$. Therefore, for predicting $Y_{t}$, $X_{t-d}$ is the optimal choice, not the freshest feature $X_{t}$. Hence, fresh data is not always the best.
@@ V-C An Optimal Scheduling Solution
The scheduling problem (34) is an infinite-horizon average-cost semi-Markov decision process (SMDP) [48, Chapter 5.6]. Define $\tau=S_{i+1}-D_{i}$ as the waiting time for sending the $(i+1)$-th feature after the $i$-th feature is delivered. The Bellman optimality equation of the SMDP (34) is
@@
It is crucial to note that a non-monotonic AoI function $p(\delta)$ often leads to a non-monotonic index function $\gamma(\delta)$. Consequently, the inverse function of $\gamma(\delta)$ may not exist and the inequality $\gamma(\Delta(t))\geq\beta_{b}$ in the threshold policy (35) cannot be equivalently rewritten as an inequality of the form $\Delta(t)\geq\alpha$. This distinction represents a significant departure from previous studies for minimizing either the AoI $\Delta(t)$ or its non-decreasing functions, e.g., [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. In these earlier works, the solutions were usually expressed as threshold policies in the form $\Delta(t)\geq\alpha$. Our pursuit of a simple threshold policy for minimizing general and potentially non-monotonic AoI functions was inspired by the restart-in-state formulation of the Gittins index [49, Chapter 2.6.4], [50].
@@ VI-F Asymptotic Optimality of the Proposed Scheduling Policy
[17, 18] Initially, we have $N$ channels and $M$ sources. Let $\bar{p}_{\pi}^{r}$ represent the expected long-term average inference error under policy $\pi$, where both the number of channels $rN$ and the number of bandits $rM$ are scaled by $r$. The policy $\pi_{\mathrm{our}}$ will be asymptotically optimal if $\bar{p}_{\pi_{\mathrm{our}}}^{r}\leq\bar{p}_{\pi}^{r}$ for all $\pi\in\Pi$ as $r$ approaches $\infty$, while maintaining a constant ratio $\alpha=\frac{rM}{rN}$.
@@
If the feature transmission times $T_{m,i}$ are bounded for all $m$ and $i$, then under the uniform global attractor condition provided in Appendix R, $\pi_{\mathrm{our}}$ is asymptotically optimal.
@@ B-4 $\alpha$-Loss
The $\alpha$-loss function is defined by $L_{\alpha}(y,Q_{Y})=\frac{\alpha}{\alpha-1}\left[1-Q_{Y}(y)^{\frac{\alpha-1}{\alpha}}\right]$ for $\alpha>0$ and $\alpha\neq 1$ [58, Eq. 14]. It becomes the log-loss function in the limit $\alpha\rightarrow 1$ and the 0-1 loss function in the limit $\alpha\rightarrow\infty$. The $L$-entropy and $L$-cross entropy associated with the $\alpha$-loss function are given by
@@ Appendix D Relationship among $L$-divergence, Bregman divergence, and $f$-divergence
Let us rewrite the $L$-entropy $H_{L}(Y)$ as $H_{L}(\mathbf{p}_{Y})$ to emphasize that it is a function of the probability vector $\mathbf{p}_{Y}$. If $H_{L}(\mathbf{p}_{Y})$ is continuously differentiable and strictly concave in $\mathbf{p}_{Y}$, then the $L$-divergence $D_{L}(P_{Y}||Q_{Y})$ can be expressed as [34, Section 3.5.4]
@@ Appendix H Proof of Lemma 2
We know from Pinsker’s inequality [42, Lemma 11.6.1] that
@@ Appendix I Proof of Theorem 1
Appendix M Proof of Theorem 4, Theorem 5, Theorem 7, and Theorem 8
@@
The scheduling problem (140) is an infinite-horizon average-cost semi-Markov decision process (SMDP) [48, Chapter 5.6]. We provide a detailed description of the components of this problem:
@@
The infinite-horizon average-cost SMDP (140) can be solved by using dynamic programming [48, 62]. Let $h:\mathbb{Z}^{+}\mapsto\mathbb{R}$ be the relative value function associated with the average-cost SMDP (140). At time $t=D_{i}$, the optimal action $(\tau_{i+1},b_{i+1})$ can be determined by solving the following Bellman optimality equation [48, P. 275]:
@@
Because $g(\mathbf{z})$ is a convex function and the set of sub-gradients of $g(\mathbf{z})$ is bounded [63, Proposition 4.2.3], we can apply the sub-gradient mean value theorem [63] along with (C), and (172) to obatin
@@
Because $Y_{t}=f(X_{t-d})$ and $X_{t}$ is a Markov chain, $Y_{t}\leftrightarrow X_{t-\delta}\leftrightarrow X_{t-(\delta-1)}$ is a Markov chain for all $0\leq\delta\leq d$. By the data processing inequality for $L$-conditional entropy [35, Lemma 12.1], one can show that for all $0\leq\delta\leq d$,
@@
Moreover, since $Y_{t}=f(X_{t-d})$ and $X_{t}$ is a Markov chain, $Y_{t}\leftrightarrow X_{t-\delta}\leftrightarrow X_{t-(\delta+1)}$ is a Markov chain for all $\delta\geq d$. By the data processing inequality [35, Lemma 12.1], one can show that for all $\delta\geq d$,
@@
Because we assume that the transmission times $T_{m,i}$ are bounded for all $m$ and $i$, we can find a $d_{\text{bound}}\in\mathbb{N}$ such that $0<T_{m,i}\leq d_{\text{bound}}$ for all $m$ and $i$. Then, the amount of time $d_{m}(t)$ that has been spent to send the current feature of source $m$ by time slot $t$ is also bounded and $d_{m}(t)\in\{0,1,\ldots,d_{\text{bound}}\}.$ Also, we observe from Figs. 1-5 that the inference error function $p_{m}(\delta)$ converges after a large AoI value. We can find an AoI $\delta_{\text{bound}}$ such that $p_{m}(\delta)=p_{m}(\delta_{\text{bound}})$ for all $\delta\geq\delta_{\text{bound}}$ and for all $m$.
@@
We define the vectors $\mathbf{V}^{m}(t)$, $\mathbf{U}^{m}(t)$, $\mathbf{v}^{m}$, and $\mathbf{u}^{m}$ to contain $(V^{m}_{\delta,d}(t),\tilde{V}^{m}_{d}(t))$, $(U^{m}_{\delta,d,\mu}(t),\tilde{U}^{m}_{d,\mu}(t))$, $(v^{m}_{\delta,d},\tilde{v}^{m}_{d})$, and $(u^{m}_{\delta,d,\mu},\tilde{u}^{m}_{d,\mu})$, respectively, for all $\delta=1,2,\ldots,\delta_{\mathrm{bound}}$, $d=0,1,\ldots,d_{\mathrm{bound}}$, and $\mu=0,1,\ldots,B_{m}$.
$\!\!\!\!\Psi_{\pi}((\mathbf{v}^{m})_{m=1}^{M}):=$
Definition 8.
@@
Policy $\pi$ in (51)-(54) can be expressed as a sequence of actions $\pi=(\mu(t))_{t=0,1,\ldots}$, with $\mu(t)=\mu_{m}(t))_{m=0}^{M}$ representing actions taken at successive time slots. Let $(\mu^{*}(t))_{t=0,1,\ldots}$ denote the optimal policy for solving (51)-(54). Given the optimal dual variable $\lambda=\lambda^{*}$, Theorems 7-8 imply that in the optimal policy for (51)-(54), class $m$ bandits choose either action $\mu_{m}(t)=0$ or $\mu_{m}(t)=b^{*}_{m}(\lambda^{*})+1$ at every time slot $t$. Therefore, the optimal state-action frequency satisfies: $u^{m*}_{\delta,d,\mu}=0$ for all $\mu\neq 0,b^{*}_{m}(\lambda^{*})+1$. Hence, for all class $m$, only two actions can occur for every bandit in class $m$. Thus, our multiple-action RMAB problem can be reduced to a binary action RMAB problem. After that, for the truncated state space $(\delta,d)\in\{1,2,\ldots,\delta_{\text{bound}}\}\times\{0,1,\ldots,d_{\text{bound}}\}$, we directly use [18, Theorem 13 and Proposition 14] to prove the asymptotic optimality of our policy. Though [18, Theorem 13 and Proposition 14] is proved for a single class of bandits, we can use the results for multiple classes of bandits because of a similar argument provided in [64, Section 5]: we argue that having multiple classes of bandits can be represented by a single class of bandits by considering a larger state space: the state of a bandit would be $(m,\delta,d)$, where $m$ is its class.

### v2

@@ IV-A Reaction Prediction with Delay
We present the following analytical example for understanding the experimental results illustrated in Figs. 2-3.
@@
Consider a causal system represented by $Y_{t}=f(X_{t-d})\in\mathcal{Y}$, where $X_{t}\in\mathcal{X}$ and $Y_{t}\in\mathcal{Y}$ are the input and output of the system, respectively, $d\geq 0$ is the delay introduced by the system, and $f(\cdot)$ is a function.
@@
Let $X_{t}\in\mathcal{X}$ and $Y_{t}=f(X_{t-d})\in\mathcal{Y}$, where $f(\cdot)$ is a function. If $X_{t}\leftrightarrow X_{t-\mu}\leftrightarrow X_{t-\mu-\nu}$ is a Markov chain for all $\mu,\nu\geq 0$, then $H_{L}(Y_{t}|X_{t-\delta})$ decreases with $\delta$ when $0\leq\delta<d$ and increases with $\delta$ when $\delta\geq d$. In addition, for any random variable $Z\in\mathcal{X}$,
$H_{L}(Y_{t}|X_{t-d})=H_{L}(Y_{t}|Y_{t})\leq\min_{Z\in\mathcal{X}}H_{L}(Y_{t}|Z).$ (29)
@@
Lemma 4 implies that the feature $X_{t-d}\in\mathcal{X}$ achieves the minimum expected loss in predicting $Y_{t}$. Therefore, for predicting $Y_{t}$, $X_{t-d}$ is the optimal choice in $\mathcal{X}$, not the freshest feature $X_{t}\in\mathcal{X}$. Hence, fresh data is not always the best.
@@ V-C An Optimal Scheduling Solution
The scheduling problem (34) is an infinite-horizon average-cost semi-Markov decision process (SMDP) [50, Chapter 5.6]. Define $\tau=S_{i+1}-D_{i}$ as the waiting time for sending the $(i+1)$-th feature after the $i$-th feature is delivered. The Bellman optimality equation of the SMDP (34) is
@@
It is crucial to note that a non-monotonic AoI function $p(\delta)$ often leads to a non-monotonic index function $\gamma(\delta)$. Consequently, the inverse function of $\gamma(\delta)$ may not exist and the inequality $\gamma(\Delta(t))\geq\beta_{b}$ in the threshold policy (35) cannot be equivalently rewritten as an inequality of the form $\Delta(t)\geq\alpha$. This distinction represents a significant departure from previous studies for minimizing either the AoI $\Delta(t)$ or its non-decreasing functions, e.g., [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. In these earlier works, the solutions were usually expressed as threshold policies in the form $\Delta(t)\geq\alpha$. Our pursuit of a simple threshold policy for minimizing general and potentially non-monotonic AoI functions was inspired by the restart-in-state formulation of the Gittins index [51, Chapter 2.6.4], [52].
@@ VI-F Asymptotic Optimality of the Proposed Scheduling Policy
[17, 18] Initially, we have $N$ channels and $M$ bandits. Let $\bar{p}_{\pi}^{r}$ represent the expected long-term average inference error under policy $\pi$, where both the number of channels $rN$ and the number of bandits $rM$ are scaled by $r$. The policy $\pi_{\mathrm{our}}$ will be asymptotically optimal if $\bar{p}_{\pi_{\mathrm{our}}}^{r}\leq\bar{p}_{\pi}^{r}$ for all $\pi\in\Pi$ as $r$ approaches $\infty$, while maintaining a constant ratio $\alpha=\frac{rM}{rN}$.
We denote by $V^{m}_{\delta,d}(t)$ the fraction of class $m$ bandits with $\Delta_{m}(t)=\delta$ and $d_{m}(t)=d$. Also, define variables $v^{m}_{\delta,d}$ for all $\delta,d$, and $m$ as follows:
$v^{m}_{\delta,d}=\limsup_{T\rightarrow\infty}\sum_{t=0}^{T-1}\frac{1}{T}\mathbb{E}[V^{m}_{\delta,d}(t)].$ (73)
If the transmission times $T_{m,i}$ are bounded for all $m$ and $i$, then we can find a $d_{\text{bound}}\in\mathbb{N}$ such that $0<T_{m,i}\leq d_{\text{bound}}$ for all $m$ and $i$. Then, the amount of time $d_{m}(t)$ that has been spent to send the current feature of source $m$ by time slot $t$ is also bounded, i.e., $d_{m}(t)\in\{0,1,\ldots,d_{\text{bound}}\}.$ Also, we observe from Figs. 1-5 that the inference error function $p_{m}(\delta)$ converges after a large AoI value. We can find an AoI $\delta_{\text{bound}}$ such that $p_{m}(\delta)=p_{m}(\delta_{\text{bound}})$ for all $\delta\geq\delta_{\text{bound}}$ and for all $m$. Therefore, we can analyze using a truncated state space $(\delta,d)\in\{1,2,\ldots,\delta_{\text{bound}}\}\times\{0,1,\ldots,d_{\text{bound}}\}$.
We further define the vectors $\mathbf{V}^{m}(t)$ and $\mathbf{v}^{m}$ to contain $V^{m}_{\delta,d}$ and $v^{m}_{\delta,d}$, respectively, for all $\delta=1,2,\ldots,\delta_{\mathrm{bound}}$, and $d=0,1,\ldots,d_{\mathrm{bound}}$. Here, we consider a truncated state space.
$\!\!\!\!\Psi_{\pi}((\mathbf{v}^{m})_{m=1}^{M})=$
Definition 6.
@@
If the feature transmission times $T_{m,i}$ are bounded for all $m$ and $i$, then under the uniform global attractor condition provided in Definition 6, $\pi_{\mathrm{our}}$ is asymptotically optimal.
@@ B-4 $\alpha$-Loss
The $\alpha$-loss function is defined by $L_{\alpha}(y,Q_{Y})=\frac{\alpha}{\alpha-1}\left[1-Q_{Y}(y)^{\frac{\alpha-1}{\alpha}}\right]$ for $\alpha>0$ and $\alpha\neq 1$ [60, Eq. 14]. It becomes the log-loss function in the limit $\alpha\rightarrow 1$ and the 0-1 loss function in the limit $\alpha\rightarrow\infty$. The $L$-entropy and $L$-cross entropy associated with the $\alpha$-loss function are given by
@@ Appendix D Relationship among $L$-divergence, Bregman divergence, and $f$-divergence
Let us rewrite the $L$-entropy $H_{L}(Y)$ as $H_{L}(\mathbf{p}_{Y})$ to emphasize that it is a function of the probability vector $\mathbf{p}_{Y}$. If $H_{L}(\mathbf{p}_{Y})$ is continuously differentiable and strictly concave in $\mathbf{p}_{Y}$, then the $L$-divergence $D_{L}(P_{Y}||Q_{Y})$ can be expressed as [37, Section 3.5.4]
@@ Appendix H Proof of Lemma 2
We know from Pinsker’s inequality [45, Lemma 11.6.1] that
@@ Appendix I Proof of Theorem 1
Appendix M Proofs of Theorem 4, Theorem 5, Theorem 7, and Theorem 8
@@
The scheduling problem (142) is an infinite-horizon average-cost semi-Markov decision process (SMDP) [50, Chapter 5.6]. We provide a detailed description of the components of this problem:
@@
The infinite-horizon average-cost SMDP (142) can be solved by using dynamic programming [50, 64]. Let $h:\mathbb{Z}^{+}\mapsto\mathbb{R}$ be the relative value function associated with the average-cost SMDP (142). At time $t=D_{i}$, the optimal action $(\tau_{i+1},b_{i+1})$ can be determined by solving the following Bellman optimality equation [50, P. 275]:
@@
Because $g(\mathbf{z})$ is a convex function and the set of sub-gradients of $g(\mathbf{z})$ is bounded [65, Proposition 4.2.3], we can apply the sub-gradient mean value theorem [65] along with (C), and (174) to obatin
@@
Because $Y_{t}=f(X_{t-d})$ and $X_{t}$ is a Markov chain, $Y_{t}\leftrightarrow X_{t-\delta}\leftrightarrow X_{t-(\delta-1)}$ is a Markov chain for all $0\leq\delta\leq d$. By the data processing inequality for $L$-conditional entropy [38, Lemma 12.1], one can show that for all $0\leq\delta\leq d$,
@@
Moreover, since $Y_{t}=f(X_{t-d})$ and $X_{t}$ is a Markov chain, $Y_{t}\leftrightarrow X_{t-\delta}\leftrightarrow X_{t-(\delta+1)}$ is a Markov chain for all $\delta\geq d$. By the data processing inequality [38, Lemma 12.1], one can show that for all $\delta\geq d$,
@@
Recall that the transmission times $T_{m,i}$ are assumed to be bounded for all $m$ and $i$. Hence, we can find a $d_{\text{bound}}\in\mathbb{N}$ such that $0<T_{m,i}\leq d_{\text{bound}}$ for all $m$ and $i$ and $d_{m}(t)\in\{0,1,\ldots,d_{\text{bound}}\}.$ Also, we can find an AoI $\delta_{\text{bound}}$ such that $p_{m}(\delta)=p_{m}(\delta_{\text{bound}})$ for all $\delta\geq\delta_{\text{bound}}$ and for all $m$.
@@
Policy $\pi$ in (51)-(54) can be expressed as a sequence of actions $\pi=(\mu(t))_{t=0,1,\ldots}$, with $\mu(t)=\mu_{m}(t))_{m=0}^{M}$ representing actions taken at successive time slots. Let $(\mu^{*}(t))_{t=0,1,\ldots}$ denote the optimal policy for solving (51)-(54). Given the optimal dual variable $\lambda=\lambda^{*}$, Theorems 7-8 imply that in the optimal policy for (51)-(54), class $m$ bandits choose either action $\mu_{m}(t)=0$ or $\mu_{m}(t)=b^{*}_{m}(\lambda^{*})+1$ at every time slot $t$. Therefore, the optimal state-action frequency satisfies: $u^{m*}_{\delta,d,\mu}=0$ for all $\mu\neq 0,b^{*}_{m}(\lambda^{*})+1$. Hence, for all class $m$, only two actions can occur for every bandit in class $m$. Thus, our multiple-action RMAB problem can be reduced to a binary action RMAB problem. After that, for the truncated state space $(\delta,d)\in\{1,2,\ldots,\delta_{\text{bound}}\}\times\{0,1,\ldots,d_{\text{bound}}\}$, we directly use [18, Theorem 13 and Proposition 14] to prove the asymptotic optimality of our policy. Though [18, Theorem 13 and Proposition 14] is proved for a single class of bandits, we can use the results for multiple classes of bandits because of a similar argument provided in [66, Section 5]: we argue that having multiple classes of bandits can be represented by a single class of bandits by considering a larger state space: the state of a bandit would be $(m,\delta,d)$, where $m$ is its class.